A body is thrown vertically upward with 55m/s g=10m/s distance travelled by the body in 6sec is

Grade:11

1 Answers

Divya
6 years ago
Meghla...hi in the question you have given; u=55m/s, g(a)=10m/s^2 and t=6s.You can use the second equation of motion i.e., s(height)=ut+1/2at^2.But as the body is thrown upwards g(acceleration) will act against the motion of the body so g will be taken as -10m/s^2. Therefore,the ans is s=u×t-1/2×a×t^2...(as g is -ve) . Which is 150 m.